# JanusVLN: Decoupling Semantics and Spatiality with Dual Implicit Memory for Vision-Language Navigation

### v1

@@ 1 Introduction
Vision-and-Language Navigation (VLN) is a foundational task in embodied AI, requiring an agent to navigate through unseen environments guided by visual inputs and natural language instructions. Recently, capitalizing on the advanced visual perception and semantic understanding capabilities of Multimodal Large Language Models (MLLMs), a new line of research (Zhang et al., 2025b; Wei et al., 2025a) has emerged. These approaches leverage vast-scale training data to adapt MLLMs into VLN models, thereby reshaping the future landscape of VLN research.
To support navigation models in conducting prolonged and effective exploration, these approaches typically only construct an explicit semantic memory. One class of methods (Zhang et al., 2025c; Zeng et al., 2024) builds a semantic cognitive map using textual descriptions for object nodes and relational edges. However, purely textual descriptions struggle to precisely convey the spatial relationships and orientation of objects, leading to the loss of crucial visual, spatial-geometric, and contextual information. Moreover, repetitive descriptions introduce substantial redundancy and noise. Another class of methods (Zhang et al., 2025b; Wang et al., 2024b) stores historical observation frames, which necessitates reprocessing the entire history of observations along with the current frame at each action prediction step, resulting in significant redundant computation. Finally, in both types of approaches, the explicit semantic memory grows exponentially as navigation time increases. This makes it exceedingly difficult for the model to extract critical information from a vast, cluttered, and fragmented memory, thereby leading to severe inefficiency.
@@
Inspired by the human brain’s hemispheric specialization for navigation, where the left hemisphere handles semantic understanding and the right manages 3D spatial cognition to form implicit representations (Gazzaniga, 1967), we propose a fundamental shift from a single, explicit memory to a dual, implicit neural memory. To this end, we introduce JanusVLN, a dual implicit memory framework for VLN that features both spatial-geometric and visual-semantic memory in Figure 1. We model these two types of memory respectively as fixed-size, compact neural memory, whose size does not grow with the trajectory length. This design is analogous to the human brain’s ability to perform efficient memorization within a finite capacity.
@@
  • Our proposed JanusVLN incrementally updates its dual implicit memory with cached KV from initial and sliding window, which eliminates the need to recompute historical frames for each new frame, substantially lowering computational and inference overhead.
@@ 2.1 Vision-language navigation with multiple visual inputs
Vision-Language Navigation (Krantz et al., 2020; Chu et al., 2025b; Li et al., 2025d), the task of guiding an embodied agent to a target location in unseen environments by following instructions, has recently garnered significant attention. Early research (Anderson et al., 2018; Sun et al., 2019) predominantly focused on discrete environments, where an agent navigates by teleporting between predefined nodes. However, these approaches (Hong et al., 2022; Zhang et al., 2024b; Zhao et al., 2025b) often exhibit poor performance when deployed on real-world robots operating in continuous 3D spaces. In contrast, more recent studies (Krantz et al., 2020; Chu et al., 2025a; Wang et al., 2025c) have concentrated on continuous environments, enabling agents to navigate freely to any collision-free location within simulators. To foster a better spatial understanding and enhance navigational capabilities, some recent works (Wang & Lee, 2025; Chen et al., 2025a; Wang et al., 2024a) have also begun to investigate monocular RGB-D vision. However, the reliance on additional, expensive hardware for this approach, which is often unavailable in many practical settings, restricts its real-world applicability. In this paper, we propose JanusVLN, a method that enhances spatial understanding using only RGB visual input, eliminating the need for any extra 3D data.
@@ 2.2 Multi-modal large language models for RGB only navigation
The recent, rapid advancement of Multi-modal Large Language Models (Bai et al., 2025; Zhang et al., 2025a; 2024d) has injected new momentum the field of Visual Language Navigation. Some approaches (Zhang et al., 2024a; Xiang et al., 2025; Zhang et al., 2025d) have begun to leverage RGB-only video models to build monocular VLN systems, aiming for enhanced generalization and practical value. However, the agents in these studies (Zhang et al., 2025b; Xie et al., 2025; Zhang et al., 2024c) typically construct only explicit semantic memory and rely solely on a single, front RGB camera, which poses significant challenges to spatial understanding and often requires extensive auxiliary data to improve performance. In this paper, we introduce JanusVLN, a VLN framework featuring a dual implicit memory system that encompasses both spatial-geometric memory and visual-semantic memory.
@@ 2.3 Spatial reasoning via vision-language models
Increasing research (Chen et al., 2024; Zeng et al., 2025; Zhao et al., 2025a) efforts have recently aimed to advance the spatial reasoning abilities of Vision-Language Models (VLMs). Previous studies (Liu et al., 2025; Leong & Wu, 2024; Wang et al., 2025a) have predominantly centered on incorporating 3D data (e.g., point clouds, depth maps) into VLMs to infuse them with explicit spatial information. However, such methods often rely on expensive auxiliary hardware, limiting their viability in practical applications. While some recent approaches (Wu et al., 2025; Ding et al., 2025b; Tang et al., 2025) leverage spatial encoders to derive spatial information directly from videos, they require the entire sequence to be re-processed upon the arrival of each new frame, leading to significant computational redundancy. JanusVLN extracts spatial-geometric features directly from video in an online, streaming fashion. This eliminates repetitive calculations and markedly lowers the inference cost.
@@ 3.1 Preliminary
Building upon traditional 3D reconstruction, recent learning-based end-to-end methods (Wang et al., 2025b; Ding et al., 2025a) employ neural networks to encode scene priors, directly predicting 3D structures from multi-view images. VGGT (Wang et al., 2025b), which is based on a transformer feed-forward architecture, comprises three key components: an encoder for extracting single-image feature, a fusion decoder for cross-frame interaction to generate geometric tokens $G_{t}\in\mathbb{R}^{\lfloor\frac{H}{p}\rfloor\times\lfloor\frac{W}{p}\rfloor\times C}$, where $p$ is the patch size, and a task-specific prediction head for 3D attributes. The reconstruction pipeline can be formulated as:
@@ 3.2 Dual implicit memory
For the implicit neural representation, we employ a hybrid cache update strategy instead of caching all historical KV. This approach mitigates the significant memory overhead and performance degradation that arise from extended navigation sequences. The strategy partitions the memory into two components. The first is a sliding window queue $M_{sliding}$ with a capacity of $n$, which stores the KV caches of the most recent $n$ frames in a first-in, first-out manner. This mechanism ensures the model focuses on the most immediate and relevant contextual information, which is critical for real-time decision-making. When this queue reaches its capacity, the oldest frame’s cache is evicted to accommodate the current frame, enabling dynamic incremental updates. The second component permanently retains the KV cache $M_{initial}$ from the initial few frames. The model exhibits sustained high attention weights towards these initial frames, which function as ”Attention Sinks” (Xiao et al., 2024; Li et al., 2025c). These sinks provide critical global anchors for the entire navigation and effectively restore performance. By integrating these two mechanisms, we construct a dynamically updated, fixed-size implicit memory that preserves an acute perception of the recent environment while maintaining a long-term memory of information.
@@
For semantic encoder and LLM, we similarly retain the KV from the initial and sliding window. Moreover, these implicit memory and tokens can be visualized to inspect the spatial and semantic information they contain.
3.3 JanusVLN architecture
@@
Additionally, Qwen2.5-VL (Bai et al., 2025) groups spatially adjacent $\times2$ patches into a single image token to reduce computational cost, yielding $S^{\prime}_{t}\in\mathbb{R}^{\lfloor\frac{H}{2p}\rfloor\times\lfloor\frac{W}{2p}\rfloor\times C}$. For 3D spatial-geometric encoder, we employ the pre-trained encoder and fusion decoder from VGGT (Wang et al., 2025b) model to interactively encode the input frame with spatial memory into spatial-geometric token $G_{t}$.
@@ 4.1 Experimental setup
Following established methods (Zhang et al., 2025b; Li et al., 2025a), we conducted experiments on two of the most recognized VLN-CE (Krantz et al., 2020) benchmark datasets: R2R-CE (Anderson et al., 2018) and RxR-CE (Ku et al., 2020). These datasets comprise trajectories collected from Matterport3D (Chang et al., 2017) scenes using the Habitat simulator (Savva et al., 2019). Consistent with prior work (Cheng et al., 2025; Li et al., 2025b), we report performance on the unseen splits using standard VLN metrics, including Navigation Error (NE), Oracle Success Rate (OS), Success Rate (SR), Success-weighted Path Length (SPL), and normalized Dynamic Time Warping (nDTW). Among these, SR and SPL are widely regarded as the primary metrics, reflecting task completion and path efficiency, respectively.
@@
We constructed JanusVLN based on Qwen2.5-VL 7B (Bai et al., 2025) and VGGT Wang et al. (2025b). The model is trained for one epoch, during which we exclusively fine-tune the LLM and the projection layer with learning rates of 2e-5 and 1e-5, respectively, while keeping the semantic and spatial encoders frozen. We set the initial and sliding window size to 8 and 48 frames. The weight for the spatial geometric features $\lambda$ is set to 0.2. For extra data, following StreamVLN (Wei et al., 2025b), we incorporated an additional 155 K trajectories from a subset of the ScaleVLN (Wang et al., 2023c), comprising approximately 9207 K image-action pairs. Furthermore, we employed the DAgger (Ross et al., 2011) algorithm to collect 14 K trajectories (approximately 1485 K image-action pairs) from the standard R2R-CE and RxR-CE datasets.
@@ 4.2 Main results
Sim2Sim [ECCV22] Krantz & Lee (2022) $\checkmark$ $\checkmark$ $\checkmark$ 6.07 52.0 43.0 36.0 -
@@
COSMO [ICCV25] (Zhang et al., 2025e) $\checkmark$ - 56.0 47.0 40.0 -
AO-Planner [AAAI25] (Chen et al., 2025b) $\checkmark$ $\checkmark$ 5.55 59.0 47.0 33.0 -
@@
g3D-LF [CVPR25] (Wang & Lee, 2025) $\checkmark$ $\checkmark$ $\checkmark$ 5.70 59.5 47.2 34.6 -
@@
NaVid-4D [ICRA25] (Liu et al., 2025) $\checkmark$ $\checkmark$ 5.99 55.7 43.8 37.1 -
NavMorph [ICCV25] (Xuan Yao & Xu, 2025) $\checkmark$ $\checkmark$ 5.75 56.9 47.9 33.2 -
NaVid [RSS24] (Zhang et al., 2024a) $\checkmark$ 5.47 49.1 37.4 35.9 $953K$
Sim2Real [CoRL24] (Wang et al., 2024c) $\checkmark$ 5.95 55.8 44.9 30.4 $0K$
StreamVLN* [arXiv25] Wei et al. (2025b) $\checkmark$ 6.05 53.8 45.5 41.6 $10033K$
@@
StreamVLN [arXiv25] Wei et al. (2025b) $\checkmark$ 4.98 64.2 56.9 51.9 $\sim 26330K$
@@
AO-Planner [AAAI25] (Chen et al., 2025b) $\checkmark$ $\checkmark$ 7.06 43.3 30.5 50.1 -
@@
NavMorph [ICCV25] (Xuan Yao & Xu, 2025) $\checkmark$ $\checkmark$ 8.85 30.8 22.8 44.2 -
Sim2Real [CoRL24] (Wang et al., 2024c) $\checkmark$ 8.79 36.7 25.5 18.1 $0K$
@@
StreamVLN [arXiv25] Wei et al. (2025b) $\checkmark$ 6.22 52.9 46.0 61.9 $\sim 26330K$
@@ Appendix A The use of large language models (LLMs)
Appendix B More ablation studies

### v2

@@ 1 Introduction
Vision-and-Language Navigation (VLN) is a foundational task in embodied AI, requiring an agent to navigate through unseen environments guided by visual inputs and natural language instructions. Recently, capitalizing on the advanced visual perception and semantic understanding capabilities of Multimodal Large Language Models (MLLMs) Lan et al. (2025a; b); Qi et al. (2026), a new line of research (Zhang et al., 2025b; Yuan et al., 2025; Liang et al., 2025; Shan et al., 2025) has emerged. These approaches leverage vast-scale training data to adapt MLLMs into VLN models, thereby reshaping the future landscape of VLN research.
To support navigation models in conducting prolonged and effective exploration, these approaches typically only construct an explicit semantic memory. One class of methods (Zhang et al., 2025c; Zeng et al., 2024; Chen et al., 2025c; Liu et al., 2025a) builds a semantic cognitive map using textual descriptions for object nodes and relational edges. However, purely textual descriptions struggle to precisely convey the spatial relationships and orientation of objects, leading to the loss of crucial visual, spatial-geometric, and contextual information. Moreover, repetitive descriptions introduce substantial redundancy and noise. Another class of methods (Cheng et al., 2025; Xiang et al., 2025; Yang et al., 2025c; Li et al., 2025a) stores historical observation frames, which necessitates reprocessing the entire history of observations along with the current frame at each action prediction step, resulting in significant redundant computation. Finally, in both types of approaches, the explicit semantic memory grows exponentially as navigation time increases. This makes it exceedingly difficult for the model to extract critical information from a vast, cluttered, and fragmented memory, thereby leading to severe inefficiency.
@@
Inspired by the human brain’s hemispheric specialization for navigation, where the left hemisphere handles semantic understanding and the right manages 3D spatial cognition to form implicit representations (Gazzaniga, 1967; Zeng et al., 2025b; Fan et al., 2025b), we propose a fundamental shift from a single, explicit memory to a dual, implicit neural memory. To this end, we introduce JanusVLN, a dual implicit memory framework for VLN that features both spatial-geometric and visual-semantic memory in Figure 1. We model these two types of memory respectively as fixed-size, compact neural memory, whose size does not grow with the trajectory length. This design is analogous to the human brain’s ability to perform efficient memorization within a finite capacity.
@@
  • We unlock the potential of spatial geometric foundation models in streaming VLN. By implementing dual-window and attention fusion mechanisms in VGGT, we efficiently update and integrate historical information incrementally.
@@ 2.1 Vision-language navigation with multiple visual inputs
Vision-Language Navigation (Krantz et al., 2020; Hao et al., 2024; Wu et al., 2025b), the task of guiding an embodied agent to a target location in unseen environments by following instructions, has recently garnered significant attention. Early research (Anderson et al., 2018; Wu et al., 2024; Zheng et al., 2025b; Zhang et al., 2025e) predominantly focused on discrete environments, where an agent navigates by teleporting between predefined nodes. However, these approaches (Hong et al., 2022; Wu and Zhang, 2025; Du et al., 2024) often exhibit poor performance when deployed on real-world robots operating in continuous 3D spaces. In contrast, more recent studies (Krantz et al., 2020; Wang et al., 2025c; Zhang et al., 2024a; 2023) have concentrated on continuous environments, enabling agents to navigate freely to any collision-free location within simulators. To foster a better spatial understanding and enhance navigational capabilities, some recent works (Wang and Lee, 2025; Xie et al., 2025b; Liu et al., 2023; Wang et al., 2025b) have also begun to investigate monocular RGB-D vision. However, the reliance on additional, expensive hardware for this approach, which is often unavailable in many practical settings, restricts its real-world applicability. In this paper, we propose JanusVLN, a method that enhances spatial understanding using only RGB visual input, eliminating the need for any supplementary 3D data.
@@ 2.2 Multi-modal large language models for RGB only navigation
The recent, rapid advancement of Multi-modal Large Language Models (Bai et al., 2025; Zeng et al., 2025c; Lin et al., 2025; Ren et al., 2025) has injected new momentum the field of Visual Language Navigation. Some approaches (Zhang et al., 2024b; Zhao et al., 2025; Yang et al., 2025a; Fan et al., 2025a) have begun to leverage RGB-only video models to build monocular VLN systems, aiming for enhanced generalization and practical value. However, the agents in these studies (Zhang et al., 2025b; Xie et al., 2025a; Cai et al., 2025; Chen et al., 2025b) typically construct only explicit semantic memory and rely solely on a single, front RGB camera, which poses significant challenges to spatial understanding and often requires extensive auxiliary data to improve performance. In this paper, we introduce JanusVLN, a VLN framework featuring a dual implicit memory system that encompasses both spatial-geometric memory and visual-semantic memory.
@@ 2.3 Spatial reasoning via vision-language models
Increasing research (Chen et al., 2024a; Zeng et al., 2025a; Zheng et al., 2025b; Zhang et al., 2025a) efforts have recently aimed to advance the spatial reasoning abilities of Vision-Language Models (VLMs). Previous studies (Chen et al., 2024b; Liu et al., 2025b; He, 2025; Ma et al., 2025) have predominantly centered on incorporating 3D data (e.g., point clouds, depth maps) into VLMs to infuse them with explicit spatial information. However, such methods often rely on expensive auxiliary hardware, limiting their viability in practical applications. While some recent approaches (Wu et al., 2025a; Zheng et al., 2025a; Jiang et al., 2025; Li et al., 2023) leverage spatial encoders to derive spatial information directly from videos, they require the entire sequence to be re-processed upon the arrival of each new frame, leading to significant computational redundancy. JanusVLN extracts spatial-geometric features directly from video in an online, streaming fashion. This eliminates repetitive calculations and markedly lowers the inference cost.
@@ 3.1 Preliminary
Building upon traditional 3D reconstruction, recent learning-based end-to-end methods (Wang et al., 2025a; Yang et al., 2025b; Liu et al., 2025c; Li and Sun, 2025) employ neural networks to encode scene priors, directly predicting 3D structures from multi-view images. VGGT (Wang et al., 2025a), which is based on a transformer feed-forward architecture, comprises three key components: an encoder for extracting single-image feature, a fusion decoder for cross-frame interaction to generate geometric tokens $G_{t}\in\mathbb{R}^{\lfloor\frac{H}{p}\rfloor\times\lfloor\frac{W}{p}\rfloor\times C}$, where $p$ is the patch size, and a task-specific prediction head for 3D attributes. The reconstruction pipeline can be formulated as:
@@ 3.2 Dual implicit memory
For the implicit neural representation, we employ a hybrid cache update strategy instead of caching all historical KV. This approach mitigates the significant memory overhead and performance degradation that arise from extended navigation sequences. The strategy partitions the memory into two components. The first is a sliding window queue $M_{sliding}$ with a capacity of $n$, which stores the KV caches of the most recent $n$ frames in a First-In, First-Out (FIFO) manner. This mechanism ensures the model focuses on the most immediate and relevant contextual information, which is critical for real-time decision-making. When this queue reaches its capacity, the oldest frame’s cache is evicted to accommodate the current frame, enabling dynamic incremental updates. The second component permanently retains the KV cache $M_{initial}$ from the initial few frames. The model exhibits sustained high attention weights towards these initial frames, which function as ”Attention Sinks” (Xiao et al., 2024; Yang et al., 2026; Lu et al., 2025a; Li et al., 2025c). These sinks provide critical global anchors for the entire navigation task and effectively restore performance. By integrating these two mechanisms, we construct a dynamically updated, fixed-size implicit memory that preserves an acute perception of the recent environment while maintaining a long-term memory of global task information.
@@
For both the semantic encoder and the LLM, namely Qwen2.5-VL, we employ standard KV Cache during the inference stage for acceleration, retaining only the key-value pairs within the initial and sliding windows. Furthermore, these implicit memories and tokens can be visualized to inspect the spatial and semantic information they encapsulate.
3.3 JanusVLN architecture
@@
Additionally, Qwen2.5-VL (Bai et al., 2025) groups spatially adjacent $\times2$ patches into a single image token to reduce computational cost, yielding $S^{\prime}_{t}\in\mathbb{R}^{\lfloor\frac{H}{2p}\rfloor\times\lfloor\frac{W}{2p}\rfloor\times C}$. For 3D spatial-geometric encoder, we employ the pre-trained encoder and fusion decoder from VGGT (Wang et al., 2025a) model to interactively encode the input frame with spatial memory into spatial-geometric token $G_{t}$.
@@ 4.1 Experimental setup
Following established methods (Zhang et al., 2025b; Li et al., 2025b; Li, 2025; Chu et al., 2025), we conducted experiments on two of the most recognized VLN-CE (Krantz et al., 2020) benchmark datasets: R2R-CE (Anderson et al., 2018) and RxR-CE (Ku et al., 2020). These datasets comprise trajectories collected from Matterport3D (Chang et al., 2017) scenes using the Habitat simulator (Savva et al., 2019). Consistent with prior work (Cheng et al., 2025; Dai et al., 2025; Yin et al., 2025; Lu et al., 2024), we report performance on the unseen splits using standard VLN metrics, including Navigation Error (NE), Oracle Success Rate (OS), Success Rate (SR), Success-weighted Path Length (SPL), and normalized Dynamic Time Warping (nDTW). Among these, SR and SPL are widely regarded as the primary metrics, reflecting task completion and path efficiency, respectively Wei et al. (2025); Lu et al. (2025c; b).
@@
We constructed JanusVLN based on Qwen2.5-VL 7B (Bai et al., 2025) and VGGT Wang et al. (2025a). The model is trained for one epoch, during which we exclusively fine-tune the LLM and the projection layer with learning rates of 2e-5 and 1e-5, respectively, while keeping the semantic and spatial encoders frozen. We set the initial and sliding window size to 8 and 48 frames. The weight for the spatial geometric features $\lambda$ is set to 0.2. For extra data, following StreamVLN (Wei et al., 2025), we incorporated an additional 155 K trajectories from a subset of the ScaleVLN (Wang et al., 2023c), comprising approximately 9207 K image-action pairs. Furthermore, we employed the DAgger (Ross et al., 2011) algorithm to collect 14 K trajectories (approximately 1485 K image-action pairs) from the standard R2R-CE and RxR-CE datasets.
@@ 4.2 Main results
Sim2Sim [ECCV22] Krantz and Lee (2022) $\checkmark$ $\checkmark$ $\checkmark$ 6.07 52.0 43.0 36.0 -
@@
COSMO [ICCV25] (Zhang et al., 2025d) $\checkmark$ - 56.0 47.0 40.0 -
AO-Planner [AAAI25] (Chen et al., 2025a) $\checkmark$ $\checkmark$ 5.55 59.0 47.0 33.0 -
@@
g3D-LF [CVPR25] (Wang and Lee, 2025) $\checkmark$ $\checkmark$ $\checkmark$ 5.70 59.5 47.2 34.6 -
@@
NaVid-4D [ICRA25] (Liu et al., 2025b) $\checkmark$ $\checkmark$ 5.99 55.7 43.8 37.1 -
NavMorph [ICCV25] (Xuan Yao and Xu, 2025) $\checkmark$ $\checkmark$ 5.75 56.9 47.9 33.2 -
NaVid [RSS24] (Zhang et al., 2024b) $\checkmark$ 5.47 49.1 37.4 35.9 $953K$
Sim2Real [CoRL24] (Wang et al., 2024) $\checkmark$ 5.95 55.8 44.9 30.4 $0K$
StreamVLN* [arXiv25] Wei et al. (2025) $\checkmark$ 6.05 53.8 45.5 41.6 $10033K$
@@
StreamVLN [arXiv25] Wei et al. (2025) $\checkmark$ 4.98 64.2 56.9 51.9 $\sim 26330K$
@@
AO-Planner [AAAI25] (Chen et al., 2025a) $\checkmark$ $\checkmark$ 7.06 43.3 30.5 50.1 -
@@
NavMorph [ICCV25] (Xuan Yao and Xu, 2025) $\checkmark$ $\checkmark$ 8.85 30.8 22.8 44.2 -
Sim2Real [CoRL24] (Wang et al., 2024) $\checkmark$ 8.79 36.7 25.5 18.1 $0K$
@@
StreamVLN [arXiv25] Wei et al. (2025) $\checkmark$ 6.22 52.9 46.0 61.9 $\sim 26330K$
@@ Appendix A The use of large language models (LLMs)
Appendix B Model structure details
Figure 5: Details of the implicit memory of the spatial geometric encoder.
In the original VGGT, frame attention and global cross-frame attention are executed alternately. In Figure 5 Our spatial encoder, in contrast, fuses information through interaction with a cache during the global attention process. Specifically, the tokens of the current frame first pass through frame attention to establish a local context. Then, during global attention, these current-frame tokens generate the Query. The final Key and Value are constructed by concatenating the historical KV cache with the newly generated KV from the current frame, which are then used to compute the attention. This alternating execution of frame attention and global attention is repeated.
Qwen2.5-VL employs the standard KV Cache mechanism typical of LLMs. Visual embeddings derived from new frame via the semantic encoder generate Queries within the language model. These Queries then compute attention against the Keys and Values of all historical tokens combined with the Keys and Values generated by the tokens of the current frame.
Appendix C More ablation studies
Figure 6: Quantitative experiments in the real world.
Real world quantitative results.
In our real-world experiments, we employed a Unitree Go2 robotic platform equipped with an Insta360 X5 camera to capture forward-facing RGB images. The JanusVLN model operates on a remote server with an A10 GPU, continuously processing RGB images and instructions, and sending the inference results back to the robot for execution. For quantitative real-world evaluation, we used 25 instructions, each repeated three times, covering both general and spatial understanding tasks. A trial is considered successful if the robot stops within 1 meter of the target. As shown in Figure 6, JanusVLN outperforms its variant without spatial memory across all scenarios. Notably, it achieves a 23.6% improvement on navigation tasks that require spatial understanding, which demonstrates the effectiveness of JanusVLN.
Table 6: Comparison on recent HM3D-OVON(Yokoyama et al., 2024b) val unseen.
[…]
VLFM [ICRA24] (Yokoyama et al., 2024a) 35.2 19.6
DAgRL+OD [IROS24] (Yokoyama et al., 2024b) 37.1 19.8
Uni-Navid [RSS25] (Zhang et al., 2025b) 39.5 19.8
MTU3D [ICCV25] (Zhu et al., 2025) 40.8 12.1
JanusVLN 44.9 31.7
Results on recent HM3D-OVON.
As shown in Table 6, we also test on the more diverse, updated HM3D-OVON (Yokoyama et al., 2024b) benchmark. Our approach JanusVLN surpasses SOTA methods by boosting the Success Rate (SR) from 40.8% to 44.9%, which showcases its strong generalization capabilities.